# Synthetic Object Compositions for Scalable and Accurate Learning in Detection, Segmentation, and Grounding

### v2

@@ Appendix B Detailed Related work
Open-Vocabulary detection and segmentation. Traditional object detection [girshick2015fast, ren2016faster] and segmentation [he2017mask, cheng2022maskedattentionmasktransformeruniversal] methods typically target predefined categories from standard benchmarks (e.g., COCO [lin2015microsoftcococommonobjects], OpenImages [Kuznetsova_2020], Object365 [9009553], LVIS [gupta2019lvisdatasetlargevocabulary]). Transformer-based architectures, such as DETR [Carion2020EndtoEndOD] and its derivatives [Zhu2020DeformableDD, Zhang2022DINODW, Lv2023DETRsBY, Huang2024DQDETRDW], have revolutionized end-to-end detection. Recent advances in vision-language models trained via contrastive language supervision [radford2021learning, Jia2021ScalingUV] have expanded detection to open-vocabulary settings [Zareian2020OpenVocabularyOD]. Approaches include distillation of pretrained vision-language knowledge into region-level representations [Gu2021OpenvocabularyOD, Zhong2021RegionCLIPRL], adapting contrastive models for detection [Minderer2022SimpleOO], integrating language-conditioned queries into DETER framework [Zang2022OpenVocabularyDW], and leveraging abundant image-level annotations to overcome detection annotation scarcity [Zhou2022DetectingTC, Lin2022LearningOA]. Building upon these approaches, scaling pre-training datasets has significantly advanced open-vocabulary models [li2021grounded, Yao2022DetCLIPDV, Yao2023DetCLIPv2SO]. GLIP [li2021grounded] combines detection with grounded language-image pretraining, introducing the ODInW benchmark [li2021grounded, li2022elevater]. Datasets like V3Det [wang2023v3detvastvocabularyvisual], with over 13,000 categories, further push open-vocabulary capabilities. MM-Grounding-DINO [Liu2023GroundingDM, zhao2024open], a combination of DINO [Zhang2022DINODW] and grounded pretraining, represents state-of-the-art performance alongside real-time solutions like Yolo-World [Cheng2024YOLOWorldRO]. Similarly, open-vocabulary segmentation (OVS) models [Zhou2021ExtractFD, ghiasi2022scalingopenvocabularyimagesegmentation, Li2022LanguagedrivenSS, Liang2022OpenVocabularySS] now incorporate language supervision, evaluated on standard semantic segmentation datasets (COCO-Stuff [Caesar2016COCOStuffTA], ADE20K [8100027, zhou2018semanticunderstandingscenesade20k], PASCAL VOC [everingham2011pascal], PASCAL Context [Mottaghi2014TheRO], and Cityscapes [Cordts2016Cityscapes]). Universal models like APE [shen2023aligningpromptinguniversalvisual] unify grounding, segmentation, and detection under diverse prompts [Zou2022GeneralizedDF, Zou2023SegmentEE].
@@
Synthetic datasets for detection, segmentation, and visual grounding. Synthetic datasets have emerged to provide large-scale annotations without manual labeling, with early examples such as SYNTHIA [Ros2016TheSD] and GTA5 [Richter2016PlayingData] offering useful segmentation labels via game-engine rendering but revealed domain gaps when applied to real-world images. Virtual KITTI 2 [Cabon2020VKITTI2] and Synscapes [Wrenninge2018Synscapes] significantly improved realism for driving scenes through photorealistic and physically-based rendering, and Hypersim [Roberts_2021_ICCV] offered extensive indoor data with dense segmentation, depth, and normal annotations. More recently, diffusion-based synthesis has produced high-resolution, photorealistic images with controllable layouts [Rombach2021HighResolutionIS, Nichol2021GLIDETP, Ramesh2022HierarchicalTI, Li2023GLIGENOG], realistic illumination variations [Kocsis2024LightItIM, zhang2025scaling], and scalability to open-vocabulary and pseudo-labeled data [He2022IsSD, Azizi2023SyntheticDF, lee2025molmoactactionreasoningmodels, lee2025molmoactactionreasoningmodels, lee2025molmoactactionreasoningmodels, Fan2023ScalingLO, zheng2025brokentokenslanguagemodel, Zhang2023AddingCC, Jiang2022PseudoQGP, Peng2023Kosmos2GM, du2025paintoutsideboxsynthesizing, gao2024generatesceneevaluatingimproving, zhang2024provision, gao2024generatesceneevaluatingimproving, zhang2024provision, gao2024generatesceneevaluatingimproving, zhang2024provision].
@@
In contrast, our framework generates over 20 million high-quality object segments across more than 47000 categories, each annotated with fine-grained metadata. We integrate a relighting-and-blending pipeline to boost photorealism and to provide precise masks, bounding boxes, and referring expressions. This strategy combines the advantages of both copy-paste and diffusion-based approaches while avoiding unrealistic artifacts and inaccurate labels. It not only overcomes the scale and category limitations of prior augmentation methods but also delivers data that is fully suitable for open-vocabulary detection, segmentation, and visual grounding in authentic real-world scenes.
@@ Appendix D Compare with Other Synthetic Method
Most synthetic datasets are treated as mere “add‑ons,” handy for augmentation yet unable to rival real‑image corpora. SOC upends that narrative. As §4 demonstrates, models trained solely on SOC already match—or surpass—the performance achieved with leading real datasets such as GRIT and V3Det. For that reason our main experiments benchmark SOC directly against real data. Here we zoom out to position SOC within the broader synthetic landscape. Table 6 and Table 7 summarise the distinctions that give SOC its edge.
@@ D.1 Comparison with Synthetic Object Segments
Library Diffusion model Category Scale (segments) Viewpoint diversity Subject 200K FLUX Less than 1000 200K SOC (ours) FLUX 47000+ 20M
@@
By contrast, SOC delivers 20M segments spanning 47K+ categories and explicitly samples multiple camera angles for every object prompt, ensuring rich viewpoint diversity. This two‑orders‑of‑magnitude increase in scale—coupled with fine‑grained control over category and viewpoint—unlocks far more varied, photorealistic composites than any prior segment library. Consistently, in Section 4.6 we show that a size‑matched subset of SOC already surpasses Subject 200K when training Mask2Former, underscoring the higher quality of our object‑segment pipeline.
@@ D.2 Comparison with Synthetic Data Pipelines
SOC is best seen as an object‑centric composition pipeline: it preserves the explicit layout control of classic copy‑paste yet discards any dependence on real host images. This shift—from augmenting a fixed corpus to generating complete scenes—unlocks four key advances:
  • Scene composition from scratch. Whereas copy‑paste can only graft segments onto existing photos, SOC builds an entire image on a blank canvas. A scripted layout engine places 5–20 objects at arbitrary positions, scales, and poses, breaking free of the viewpoint and co‑occurrence biases of real corpora.
  • 20 M open‑vocabulary object segments pool. A library of 20 M high‑resolution segments covering 47 K categories lifts the ceiling imposed by LVIS/COCO copy‑paste pipelines (~1.3 K classes).
  • Photorealistic harmonisation. Diffusion‑based global relighting aligns illumination and colour across foregrounds and background, while a mask‑aware blending pass restores edge detail, erasing the seam artefacts and lighting clashes of naïve copy‑paste.
  • Detector‑free ground truth. Every object is placed by script, so its mask, bounding box, and referring phrases are exact by construction; no detector or VLM is invoked, avoiding the pseudo‑label noise that hampers diffusion‑plus‑detector pipelines.
The outcome is a dataset that matches simulators in annotation fidelity, rivals diffusion images in photorealism, and, through open‑vocabulary, layout‑controlled synthesis, vastly outstrips prior copy‑paste methods in diversity and scale. These qualities underpin the improvements reported in §4.1–§4.5, where SOC not only complements but often surpasses real‑image training data—a stark contrast to existing synthetic datasets that are commonly treated as mere complements to real data.
@@ D.3 Visual Realism: FID Score Analysis
However, we emphasize that SOC optimizes for different objectives than photorealism alone: annotation accuracy (pixel-perfect masks), compositional diversity (40K+ categories with unlimited combinations), and training effectiveness. Our design choices—mask-weighted blending that preserves object attributes over seamless boundaries, and randomized layouts that maximize diversity over natural scene statistics—deliberately prioritize these objectives.
The results validate this approach: despite the FID scores still being relatively high in absolute terms, models trained on SOC outperform those trained on photorealistic real datasets GRIT (+3.8 AP on LVIS) and V3Det (+0.8 AP on LVIS), as demonstrated in Section 4.1. This demonstrates that for synthetic training data, compositional diversity and annotation accuracy matter more than human-perceived realism as measured by FID.
@@ E.1 Details of Mask-Area-Weighted Blending Algorithm
    $r_{i}$ $=\frac{\operatorname{area}(M_{i})-A_{\min}}{A_{\max}-A_{\min}},$ $r_{i}\in[0,1],$ (3)
    $\alpha_{i}$ $=\alpha_{\min}+(\alpha_{\max}-\alpha_{\min})\,\sigma\!\left(s\,(r_{i}-\tfrac{1}{2})\right).$ (4)
@@
    $L_{\text{out}}$ $=\alpha_{i}L_{O}+(1-\alpha_{i})L_{R},$ (5)
    $\mathbf{c}_{\text{out}}$ $=(1-\beta_{i})\,\mathbf{c}_{O}+\beta_{i}\,\mathbf{c}_{R},\quad\mathbf{c}=(a,b),\ \ \beta_{i}\ll 1,$ (6)
@@
E.2 Prompts used in SOC

### v3

@@ Appendix B Detailed Related work
Open-Vocabulary detection and segmentation. Traditional object detection [girshick2015fast, ren2016faster] and segmentation [he2017mask, cheng2022maskedattentionmasktransformeruniversal] methods typically target predefined categories from standard benchmarks (e.g., COCO [lin2015microsoftcococommonobjects], OpenImages [Kuznetsova_2020], Object365 [9009553], LVIS [gupta2019lvisdatasetlargevocabulary]). Transformer-based architectures, such as DETR [Carion2020EndtoEndOD] and its derivatives [Zhu2020DeformableDD, Zhang2022DINODW, Lv2023DETRsBY, Huang2024DQDETRDW], have revolutionized end-to-end detection. Recent advances in vision-language models trained via contrastive language supervision [radford2021learning, Jia2021ScalingUV] have expanded detection to open-vocabulary settings [Zareian2020OpenVocabularyOD]. Approaches include distillation of pretrained vision-language knowledge into region-level representations [Gu2021OpenvocabularyOD, Zhong2021RegionCLIPRL], adapting contrastive models for detection [Minderer2022SimpleOO], integrating language-conditioned queries into DETR framework [Zang2022OpenVocabularyDW], and leveraging abundant image-level annotations to overcome detection annotation scarcity [Zhou2022DetectingTC, Lin2022LearningOA]. Building upon these approaches, scaling pre-training datasets has significantly advanced open-vocabulary models [li2021grounded, Yao2022DetCLIPDV, Yao2023DetCLIPv2SO]. GLIP [li2021grounded] combines detection with grounded language-image pretraining, introducing the ODInW benchmark [li2021grounded, li2022elevater]. Datasets like V3Det [wang2023v3detvastvocabularyvisual], with over 13,000 categories, further push open-vocabulary capabilities. MM-Grounding-DINO [Liu2023GroundingDM, zhao2024open], a combination of DINO [Zhang2022DINODW] and grounded pretraining, represents state-of-the-art performance alongside real-time solutions like Yolo-World [Cheng2024YOLOWorldRO]. Similarly, open-vocabulary segmentation (OVS) models [Zhou2021ExtractFD, ghiasi2022scalingopenvocabularyimagesegmentation, Li2022LanguagedrivenSS, Liang2022OpenVocabularySS] now incorporate language supervision, evaluated on standard semantic segmentation datasets (COCO-Stuff [Caesar2016COCOStuffTA], ADE20K [8100027, zhou2018semanticunderstandingscenesade20k], PASCAL VOC [everingham2011pascal], PASCAL Context [Mottaghi2014TheRO], and Cityscapes [Cordts2016Cityscapes]). Universal models like APE [shen2023aligningpromptinguniversalvisual] unify grounding, segmentation, and detection under diverse prompts [Zou2022GeneralizedDF, Zou2023SegmentEE].
@@
Synthetic datasets for detection, segmentation, and visual grounding. Synthetic datasets have emerged to provide large-scale annotations without manual labeling, with early examples such as SYNTHIA [Ros2016TheSD] and GTA5 [Richter2016PlayingData] offering useful segmentation labels via game-engine rendering but revealed domain gaps when applied to real-world images. Virtual KITTI 2 [Cabon2020VKITTI2] and Synscapes [Wrenninge2018Synscapes] significantly improved realism for driving scenes through photorealistic and physically-based rendering, and Hypersim [Roberts_2021_ICCV] offered extensive indoor data with dense segmentation, depth, and normal annotations. More recently, diffusion-based synthesis has produced high-resolution, photorealistic images with controllable layouts [Rombach2021HighResolutionIS, Nichol2021GLIDETP, Ramesh2022HierarchicalTI, Li2023GLIGENOG], realistic illumination variations [Kocsis2024LightItIM, zhang2025scaling], and scalability to open-vocabulary and pseudo-labeled data [He2022IsSD, Azizi2023SyntheticDF, lee2025molmoactactionreasoningmodels, Fan2023ScalingLO, zheng2025brokentokenslanguagemodel, Zhang2023AddingCC, Jiang2022PseudoQGP, Peng2023Kosmos2GM, du2025paintoutsideboxsynthesizing, gao2024generatesceneevaluatingimproving, zhang2024provision].
@@
In contrast, our framework generates over 20 million high-quality object segments across more than 46K categories, each annotated with fine-grained metadata. We integrate generative harmonization (IC-Light for background inpainting and global relighting) and mask-area-weighted blending to boost photorealism and to provide precise masks, bounding boxes, and referring expressions. This strategy combines the advantages of both copy-paste and diffusion-based approaches while avoiding unrealistic artifacts and inaccurate labels. It not only overcomes the scale and category limitations of prior augmentation methods but also delivers data that is fully suitable for open-vocabulary detection, segmentation, and visual grounding in authentic real-world scenes.
@@ Appendix D Compare with Other Synthetic Method
We position SOC within the broader synthetic landscape through qualitative comparisons in Table 6 and Table 8.
@@ D.1 Comparison with Synthetic Object Segments
Library Diffusion model Category Scale (segments) Viewpoint diversity Subject 200K FLUX Less than 1000 200K SOC (ours) FLUX 46K+ 20M
@@
By contrast, SOC delivers 20M segments spanning 46K+ categories and explicitly samples multiple camera angles for every object prompt, ensuring rich viewpoint diversity. This two‑orders‑of‑magnitude increase in scale—coupled with fine‑grained control over category and viewpoint—unlocks far more varied, photorealistic composites than any prior segment library.
Quantitative comparison with Subject200K.
To validate the quality advantage of our synthetic object segments, we conduct a controlled experiment on COCO instance segmentation using Mask2Former trained on 10K synthetic images. As shown in Table 7, when combining real segments with synthetic object segments, SOC achieves 12.79 AP, outperforming Subject200K (12.06 AP) by +0.73 AP (+6.1% relative improvement). This demonstrates that our object segment pipeline—with its broader category coverage (46K+ vs. <1K) and multi-view diversity—produces higher-quality segments that lead to better downstream performance.
Table 7: Quantitative comparison with Subject200K on COCO instance segmentation (Mask2Former, zero-shot). All models are trained on 10K synthetic images composed from real segments + synthetic object segments.
[…]
Real segments only 7.03 0.55 3.15 7.22
Real + Subject200K 12.06 1.43 6.53 14.57
Real + SOC (ours) 12.79 1.58 7.71 15.74
Improvement +0.73 +0.15 +1.18 +1.17
@@ D.2 Comparison with Synthetic Data Pipelines
SOC is best seen as an object‑centric composition pipeline. The detailed comparisons are shown in Table 6 and Table 8.
The outcome is a dataset that matches simulators in annotation fidelity, rivals diffusion images in photorealism, and, through open‑vocabulary, layout‑controlled synthesis, vastly outstrips prior copy‑paste methods in diversity and scale. These qualities underpin the improvements reported in Sec. 4.14.5, where SOC demonstrates strong performance against both synthetic-based methods and real-image training data (e.g., GRIT, V3Det)—a stark contrast to existing synthetic datasets that are commonly treated as mere complements to real data.
@@ D.3 Visual Realism: FID Score Analysis
We note that SOC optimizes for annotation accuracy (pixel-perfect masks), compositional diversity (46K+ categories), and training effectiveness rather than photorealism alone. The strong downstream performance against both synthetic-based and non-synthetic-based methods (Sec. 4.1) validates this design choice.
@@ E.1 Details of Mask-Area-Weighted Blending Algorithm
    $r_{i}$ $=\frac{\operatorname{area}(M_{i})-A_{\min}}{A_{\max}-A_{\min}},$ $r_{i}\in[0,1],$
    $\alpha_{i}$ $=\alpha_{\min}+(\alpha_{\max}-\alpha_{\min})\,\sigma\!\left(s\,(r_{i}-\tfrac{1}{2})\right).$
@@
    $L_{\text{out}}$ $=\alpha_{i}L_{O}+(1-\alpha_{i})L_{R},$
    $\mathbf{c}_{\text{out}}$ $=(1-\beta_{i})\,\mathbf{c}_{O}+\beta_{i}\,\mathbf{c}_{R},\quad\mathbf{c}=(a,b),\ \ \beta_{i}\ll 1,$
@@
E.2 Details of 3D Geometric Layout Augmentation
Robustness-first design: Why photorealistic layouts are not our goal.
A central design principle of SOC is that photorealistic spatial layouts are neither necessary nor desirable for training robust vision models. Real-world photographs contain strong statistical regularities (e.g., cars appear large and near image bottoms, small objects cluster on surfaces) that models can exploit as shortcuts. Our 3D geometric layout augmentation deliberately breaks these correlations by sampling object depth, size, and position independently of category, creating compositions that may appear “unnatural” but force models to learn view-invariant, category-robust representations.
Generative harmonization plays a complementary role: it eliminates low-level copy-paste artifacts (lighting inconsistencies, boundary discontinuities) that would provide trivial cues for distinguishing synthetic data, without constraining spatial layouts to match photographic distributions. This separation is intentional—harmonization ensures sufficient visual coherence to avoid shortcut learning from obvious artifacts, while our layout strategy ensures sufficient diversity to avoid shortcut learning from statistical priors.
The effectiveness of this approach is demonstrated by our results (Sec. 4.14.5), where SOC-trained models outperform both synthetic-based methods and models trained on real images (GRIT, V3Det). These results suggest that photorealistic layouts may actually be detrimental to robustness, as they reintroduce the pictorial biases that limit generalization.
3D scene modeling with category-independent sampling.
Our 3D geometric layout augmentation strategy models each composite image as a 3D scene where depth and spatial position are sampled independently of object category. This ensures objects of the same category appear at diverse depths, sizes, and positions, preventing category-specific pictorial patterns. For each image, we sample 5-20 object segments (matching COCO/SA-1B distributions) using balanced category sampling to avoid bias.
Each object category $c$ has a commonsense physical size range (e.g., cars: 4-5m, cups: 10-20cm) generated by Qwen2.5-32B. The complete pipeline is:
  1. 1.
    Sample camera focal length: $f\sim\mathcal{U}(f_{\min},f_{\max})$
  2. 2.
    Define maximum depth: $D_{\max}=\alpha\cdot f$ (where $\alpha$ is a scaling constant in meters/pixel)
  3. 3.
    Define depth ranges (in meters):
    • Close: $0.1D_{\max},0.3D_{\max}]$
    • Middle: $0.3D_{\max},0.6D_{\max}]$
    • Far: $0.6D_{\max},D_{\max}]$
  4. 4.
    For each object segment $i$ of category $c_{i}$:
    • Sample physical size: $S_{i}\sim\mathcal{N}(\mu_{c_{i}},\sigma_{c_{i}})$
    • Sample depth $d_{i}$ from one of the three ranges, following COCO/SA-1B distribution (40% close, 35% middle, 25% far)
    • Sample 3D position: $(X_{i},Y_{i})\sim\mathcal{U}(X_{\min},X_{\max})\times\mathcal{U}(Y_{\min},Y_{\max})$ (in meters)
    • Project to 2D via perspective projection:
      $x_{i}=f\cdot\frac{X_{i}}{d_{i}},\quad y_{i}=f\cdot\frac{Y_{i}}{d_{i}},\quad s_{i}=f\cdot\frac{S_{i}}{d_{i}}$
      where $(x_{i},y_{i})$ is the 2D center position and $s_{i}$ is the apparent size in pixels
  5. 5.
    Enforce constraints: If an object’s apparent size is too small/large, or if it completely occludes another object ($\text{IoU}(M_{i},M_{j})\geq 0.9$), resample its 3D position and depth
This approach ensures that object scale is determined by 3D geometry (depth + physical size) rather than category, breaking spurious correlations like “cars appear large and near the bottom.”
E.3 Details of Camera Configuration Augmentation
After composing and relighting the scene, we apply camera configuration augmentation to simulate diverse camera intrinsics and viewing conditions. Each augmentation (random zoom and depth-of-field blur) is applied independently with 30% probability.
Random zoom (scaling and cropping).
Starting from the focal length $f$ sampled during layout generation, we apply random scaling with factor $s\sim\mathcal{U}(1.0,4.0)$ followed by random cropping to simulate camera zoom in. For a composite image $I$ of size $H\times W$:
  1. 1.
    Resize to $sH\times sW$ (modifying the focal length to $f^{\prime}=s\cdot f$)
  2. 2.
    Randomly crop back to $H\times W$
This operation ensures that object scale is not a reliable cue for category recognition.
Depth-of-field blur.
To simulate realistic depth-of-field effects controlled by aperture size, we apply selective Gaussian blur based on object depth:
  1. 1.
    Generate a depth map for the composed image: We first use Depth Anything V2 to predict relative depth for the entire scene, then scale it such that the predicted depth of the farthest object matches its sampled depth $d_{\max}$ from the 3D scene modeling. Finally, we replace each object region with its exact sampled depth $d_{i}$. This approach ensures physical consistency with the layout while allowing natural depth variation in the background
  2. 2.
    Randomly sample a focal plane depth $d_{\text{focal}}$ from the scene’s depth distribution
  3. 3.
    Sample an f-number $N\sim\mathcal{U}(1.4,16)$ representing the aperture size (smaller f-numbers = larger apertures = shallower depth-of-field)
  4. 4.
    Compute blur kernel size for each pixel at depth $d$ via the circle of confusion formula:
    $\sigma(d)=\frac{f^{2}}{N\cdot d_{\text{focal}}}\cdot\frac{|d-d_{\text{focal}}|}{d}$
    where $f$ is the focal length sampled during layout generation and $d$ is the absolute depth in meters from the depth map
Objects near the focal plane remain sharp ($\sigma\approx 0$), while those farther away are progressively blurred. Smaller f-numbers (e.g., f/1.4) produce strong background blur mimicking portrait photography, while larger f-numbers (e.g., f/16) keep most objects in focus, simulating landscape photography.
These camera configuration augmentations, combined with our 3D geometric layout augmentation strategy, create a rich distribution of visual configurations that force models to learn robust, view-invariant representations.
Figure 8: Examples of random zoom augmentation. Starting from the composed image (left), we apply random scaling with factor $s\sim\mathcal{U}(1.0,4.0)$ followed by random cropping (right). This simulates camera zoom in and ensures object scale is not a reliable cue for category recognition.
Figure 9: Examples of depth-of-field blur augmentation. Starting from the composed image (left), we apply selective Gaussian blur based on depth (right). Objects near the focal plane remain sharp while those farther away are progressively blurred, simulating realistic camera aperture effects with varying f-numbers.
E.4 Prompts used in SOC
We provide our prompts used in belowing images.
@@ Appendix H Quality of Synthetic Segments
Appendix I Gallery of Object Segments
Appendix J Gallery of Training Data (Before Camera Configuration Augmentation)
Note: Camera configuration augmentation (random zoom and depth-of-field blur) is applied on-the-fly during training. The images shown here are before this augmentation step. See pages 1- of figures/training_data_example.pdf